Tin (II) fluoride, also known as stannous fluoride, is added to some dental products to help prevent
cavities. How many moles of tin (II) fluoride can be made from 55.0 g of hydrogen fluoride, HF, if there
is plenty of tin?
1
Expert's answer
2020-02-11T05:38:14-0500
Sn+2HF=>SnF2+H2
Molar mass of HF=1+19=20g/mol
Moles of HF=55g/(20g/mol)=55/20=2.75mol
From 2 moles of HF we can get 1 mol of SnF2
according to the equation
We have 2.75 moles of HF, so 
moles of SnCl2=2.75/2=1.375 moles
Answer:1.375 moles